OpenStudy (anonymous):

I have a algebra 2 question I have no idea how to solve: You can type 1 page in 20 minutes. Your friend can type 1 page in 25 minutes. How long will it take you and your friend to type 60 pages if you work together?

OpenStudy (lgbasallote):

you use the formula \[\huge \frac tx + \frac ty = 1\] t is the total time it takes working together x is the time it takes an individual to finsih alone y is the time it takes for another individual to finish alone first you need to get how long you can type 60 pages alone you can typ 1 page in 20 minutes so you can time 60 pages in 1200 minutes.. make sense? your friend can type 1 page in 25 minutes so he can type 60 pages in 1500 minutes so \[\huge \frac{t}{1200} + \frac{t}{1500} = 1\] solve for t does that help?

OpenStudy (lgbasallote):

or do you need more help?

OpenStudy (anonymous):

THANK YOU SO MUCH!

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

how would you go about solving for t with the denominators different?

OpenStudy (lgbasallote):

are you asking me or are you trying to give a hint?

OpenStudy (anonymous):

you...

OpenStudy (lgbasallote):

you multiply all terms by LCM for example you have \[\huge \frac t2 + \frac t3 = 1\] you multiply all terms by the LCM which in this case is 6 so \[\huge \implies 3t + 2t = 6\] add like terms \[\huge 5t = 6\] divide by 5 \[\huge t = \frac 65 \; \text{hours}\]

OpenStudy (anonymous):

okay need more help

OpenStudy (anonymous):

i accidentally added the two different denominators together whoops... what comes after that.

OpenStudy (anonymous):

oh thanks.

OpenStudy (lgbasallote):

like i said you have \[\huge \frac{t}{1200} + \frac{t}{1500} = 1\] the next step is to multiply all terms by the LCM hint: the LCM for this one is 6,000

OpenStudy (lgbasallote):

^btw.. these are minutes

OpenStudy (anonymous):

okay so far i have \[\frac{ 54 }{ 600 }\]

OpenStudy (anonymous):

1200t+1500t = 6000 2700t = 6000 t = 6000/2700 or 2.22

OpenStudy (lgbasallote):

how did you get 54/600 again?

OpenStudy (anonymous):

I crossed out the zeros in the fraction... it was \[\frac{ 540 }{ 6000 } \] but then i simplifyed it.

OpenStudy (anonymous):

is this the only equation you can use to solve this type of problem or is this the one you know?

OpenStudy (anonymous):

by equation i mean formula*

OpenStudy (lgbasallote):

it actually comes from the formula \[\frac 1x + \frac 1y = \frac 1t\]

OpenStudy (anonymous):

only one LOL. yeah i have a test tomorrow. I believe there is only one question like this. thank goodness

OpenStudy (lgbasallote):

you can also use \[t = \frac{ab}{a+b}\]

OpenStudy (lgbasallote):

by a and b i mean x and y respectively

OpenStudy (anonymous):

when i first saw this question i thought system of equations. but i've only done Algebra I so far and will start Alegbra II this year so i haven't learned this.

OpenStudy (lgbasallote):

anyway you ahve \[\huge \frac{1}{1200} + \frac{1}{1500} = 1\] multiply all terms by 6,000 \[\huge \implies 5t + 4t = 6,000\] combine like terms \[\implies 9t = 6000\] divide both sides by 9 \[\huge\implies t = 666.6\overline 6\]

OpenStudy (lgbasallote):

^minutes

OpenStudy (anonymous):

Thank you, that makes sense. :)

OpenStudy (lgbasallote):

welcome